OpenStudy (anonymous):

what does 2c + 1 =-31

jimthompson5910 (jim_thompson5910):

2c + 1 =-31 2c =-31-1 2c =-32 c = -32/2 c = -16

OpenStudy (anonymous):

c=16

OpenStudy (anonymous):

THANKS